OpenStudy (anonymous):

In a right triangle, the measures of the two smaller angles are (3X+6) and (2X-1) . Find the measure of the smallest angle in the triangle.

OpenStudy (compassionate):

\(\Huge{\color{purple}{\textbf{W}} \color{orange}{\cal{E}} \color{green}{\mathbb{L}} \color{blue}{\mathsf{C}} \color{maroon}{\rm{O}} \color{red}{\tt{M}} \color{gold}{\tt{E}} \space \color{orchid}{\mathbf{T}} \color{Navy}{\mathsf{O}} \space \color{OrangeRed}{\boldsymbol{O}} \color{Olive}{\mathbf{P}} \color{Lime}{\textbf{E}} \color{DarkOrchid}{\mathsf{N}} \color{Tan}{\mathtt{S}} \color{magenta}{\mathbb{T}} \color{goldenrod}{\mathsf{U}} \color{ForestGreen}{\textbf{D}} \color{Salmon}{\mathsf{Y}} \ddot \smile }\) Hi, You seem to have an incomplete answer. Please revise it, and give me your numbers. OpenStudy Ambassador; Compassionate.

OpenStudy (campbell_st):

well since the triangle is right angled you know that the given angles sum to 90........(angle sum of a triangle 180, 1 angle = 90) so you can say 3x + 6 + 2x - 1 = 90 or 5x - 5 = 90 now solve for x. the smaller angle will be 2x - 1 so substitute your answer for x and you'll have the measure of the smaller angle. hope this helps

OpenStudy (compassionate):

@campbell_st , my computer won't load the numbers he posted. Where are you getting that? Or is it just me? e.e

OpenStudy (campbell_st):

its the question you're responding to.

OpenStudy (compassionate):

His numbers don't load for me, I guess. They're just squares. I use a MAC one time and this happens. Thanks Obama.

OpenStudy (campbell_st):

here is the question In a right triangle, the measures of the two smaller angles are (3X+6) and (2X-1) . Find the measure of the smallest angle in the triangle.

OpenStudy (compassionate):

There's just square blocks were the numbers should be. My computer is racist.

OpenStudy (fellowroot):

its 33 degrees

OpenStudy (campbell_st):

thats correct @fellowroot, but the idea of Open Study is to help people with their learning rather than just giving answers...

OpenStudy (campbell_st):

well the question is In a right triangle the smaller angles are 3x+ 6 and 2x -1 . find the measure of the smallest angle.

OpenStudy (compassionate):

Now I see it. My computer must be prejudice against parenthesis.

OpenStudy (campbell_st):

lol..... ok... they do tend to stick together

OpenStudy (fellowroot):

giving out the answer was only there so they can check their work. sorry

OpenStudy (campbell_st):

its fine @fellowroot... just allow the asker to respond... enjoy helping people here

OpenStudy (anonymous):

(3x+6)+(2x-1)+90=180 4x=90-5 x=85/4 x=21.25

OpenStudy (anonymous):

put x in both values and u get the smallest angle